On the complementation of C(E, F) in L(E, F) for non-archimedean fields

N. De Grande–De Kimpe, Brussel, Belgium and C. Perez–Garcia^{∗}, Santander, Spain

Keywords: p-adic functional analysis, compactoid sets, compact operators AMS Subject classification: 46S10

Abstract: For non-archimedean Banach spaces E, F, the following question is studied. “If the space C(E, F) (the compact linear maps) is complemented in L(E, F) (the continuous linear maps), does it imply that C(E, F) = L(E, F)?”.